### v1

# LLM-Barber: Block-Aware Rebuilder for Sparsity Mask in One-Shot for Large Language Models

Abstract
Large language models (LLMs) have grown significantly in scale, leading to a critical need for efficient model pruning techniques. Existing post-training pruning techniques primarily focus on measuring weight importance on converged dense models to determine salient weights to retain. However, they often overlook the changes in weight importance during the pruning process, which can lead to performance degradation in the pruned models. To address this issue, we present LLM-Barber (Block-Aware Rebuilder for Sparsity Mask in One-Shot), a novel one-shot pruning framework that rebuilds the sparsity mask of pruned models without any retraining or weight reconstruction. LLM-Barber incorporates block-aware error optimization across Self-Attention and MLP blocks, ensuring global performance optimization. Inspired by the recent discovery of prominent outliers in LLMs, LLM-Barber introduces an innovative pruning metric that identifies weight importance using weights multiplied by gradients. Our experiments show that LLM-Barber can efficiently prune models like LLaMA and OPT families with 7B to 13B parameters on a single A100 GPU in just 30 minutes, achieving state-of-the-art results in both perplexity and zero-shot performance across various language benchmarks. Code is available at https://github.com/YupengSu/LLM-Barber.
Introduction
Large language models (LLMs) have become a cornerstone in natural language processing (NLP) due to their impressive performance on various tasks. However, as these models increase in size and complexity, their deployment poses significant challenges due to extensive computational and storage demands. For instance, models such as GPT-175B (Brown et al. 2020), with 175 billion parameters, require vast resources, making it impractical for many applications. Therefore, efficient model compression strategies are crucial for deploying these powerful models in practical applications.
Model compression techniques commonly employ quantization and pruning to enhance model efficiency. Quantization reduces the precision of model parameters, while pruning removes less critical parameters. Traditional pruning methods, such as magnitude-based pruning (Han et al. 2015), directly trim weights based on their absolute values. While effective for smaller models, these methods often struggle with large-scale LLMs, resulting in suboptimal sparsity due to their inability to capture the complex interactions and importance of weights in such massive scalability. To address these limitations, recent post-training pruning techniques such as SparseGPT and Wanda have emerged.
Current pruning methods face two major challenges. First, as depicted in Figure 1(a), traditional layer-aware pruning methods focus on individual layers and neglect inter-layer dependencies within, leading to higher error accumulation (represented by blue arrows). In contrast, block-aware pruning, by considering groups of layers, captures inter-layer interactions to reduce error accumulation (represented by orange arrows). Second, as shown in Figure 1(b), conventional methods typically build the pruning mask once, ignoring the changes of weight significance in post-pruning stage. This oversight can lead to improper identification of salient weights, resulting in performance degradation.
Figure 1: The benefits of integrating LLM-Barber into the pruning process: (a) Transition from the layer-aware to block-aware error accumulation to achieve an optimized global solution. (b) Rebuilding sparsity mask using a novel pruning metric based on weights multiplied by gradients.
To address these limitations, we propose LLM-Barber, a novel and straightforward approach designed to rebuild sparsity mask of pruned networks without requiring for retraining or weight reconstruction. Firstly, unlike layer-aware methods that are confined to local optimization and thus prone to significant error accumulation, as depicted by the blue arrows in Figure 1(a), LLM-Barber integrates pruning across both Self-Attention and MLP block. This approach mitigates error accumulation, as evidenced by the lighter orange arrows, facilitating global optimization and improved model performance. Secondly, LLM-Barber identifies weights that, although initially non-salient without a sparsity mask, gain significance in post-pruning. As shown in Figure 1(b), varying color shades represent the relative importance scores of different weights. LLM-Barber accurately identifies and rebuilds masks for high-score (deeply shaded) weights while pruning newly identified low-score (lightly shaded) weights. Thirdly, LLM-Barber utilizes a first-order Taylor series as pruning metric, which leverages weight-gradient multiplication for precise rebuilding of the sparsity mask, thereby enabling accurate pruning decisions and significantly reducing computational complexity compared to methods using second-order information. Finally, our one-shot approach surpasses iterative fine-tuning in efficiency while maintaining comparable accuracy. By rebuilding sparsity masks of pruned model at once, LLM-Barber offers a faster and more precise pruning solution. We validate the effectiveness of LLM-Barber through extensive experiments on the LLaMA and OPT families. Our method consistently outperforms existing post-training pruning techniques in both perplexity and zero-shot performance, establishing LLM-Barber as a state-of-the-art approach in this domain. To sum up, the key contributions are fourfold:
  • Block-Aware Global Optimization: We are among the first to introduce a block-aware reconstruction problem that integrates sparsity across the Self-Attention and MLP blocks, achieving global optimization in pruning.
  • Rebuilding Sensitive Regions: We identify non-salient weights that restore significance in post-pruning stage and rebuild sparsity mask to retain these weights, while simultaneously adjusting the mask to prune newly non-salient weights. This targeted rebuilding enhances overall model performance by optimally reallocating sparsity.
  • Innovative Pruning Metric: We propose an innovative pruning metric based on the product of weights and gradients, leveraging first-order Taylor series for importance evaluation to reduce computational complexity compared to second-order Hessian-based approaches.
  • Versatility and Efficiency: LLM-Barber demonstrates its efficiency and effectiveness across various pruning techniques, consistently achieving state-of-the-art performance in perplexity and zero-shot tasks, thereby establishing new yardsticks in LLM post-training pruning.
Related Work
LLMs Pruning and Sparsity
Network pruning reduces deep neural networks by removing unnecessary weights. For LLMs, pruning methods are divided into parameter efficient fine-tuning and post-training approaches. Parameter Efficient Fine-tuning (PEFT) begins with a initialized sparse network and refines it through iterative processes (Liu et al. 2019). LoRA (Hu et al. 2021) adapts pre-trained models to specific tasks or domains by injecting trainable rank decomposition matrices. Dynamic Spase No Training (Zhang et al. 2024) minimizes the reconstruction error by iteratively pruning and growing weights. However, fine-tuning requires ample data and often leads to performance decline. Recent studies advance toward one-shot post-training pruning, showing substantial improvements. Post-training pruning removes weights from a pre-trained model. SparseGPT (Frantar and Alistarh 2023) uses Hessian-based metrics and subsequent residual weight updates, while Wanda (Sun et al. 2023) introduces a first-order pruning metric using weight-activation products. LLM-Barber employs a block-aware reconstruction approach and rebuilds masks with a novel pruning metric.
Model Compression Strategy
Compression is key to reducing the memory and computational demands of model. This work highlights the block-aware compression strategy over traditional layer-aware approaches. Layer-aware compression strategy began with Optimal Brain Damage (LeCun, Denker, and Solla 1989) and Optimal Brain Surgeon (Hassibi, Stork, and Wolff 1993). Building on these foundations, recent works like GPTQ (Frantar et al. 2023) further enhance layer-aware quantization using second-order information. Block-aware compression strategy generally offer better accuracy recovery in pruned models compared to layer-aware methods. For instance, APTQ (Guan et al. 2024) applies global quantization to attention mechanisms, enhancing model robustness, while BESA (Xu et al. 2024) uses block-wise sparsity allocation. Our method leverages block-aware pruning to optimize global performance across Self-Attention and MLP blocks, effectively balancing efficiency and accuracy.
Block-Aware Rebuilder for Sparsity Mask
Figure 2: The workflow of LLM-Barber. (a) illustrates the process of block-aware reconstruction error calculation and gradient back-propagation for each linear weight. (b) shows pruning metric computation and procedure for rebuilding sparsity mask.
Preliminaries
LLM pruning removes weights from dense networks to minimize output discrepancies, which is computationally intensive cross large-scale models, leading to address a layer-aware reconstruction problem (Hassibi, Stork, and Wolff 1993). This section reviews and reanalyses layer-aware reconstruction error and Taylor expansion at dense networks.
Layer-aware Reconstruction Error. For linear projection layer weight $\mathbf{W}$ of shape ${(C_{\text{out}},C_{\text{in}})}$, where $C_{\text{out}}$, $C_{\text{in}}$ indicates the output and input channels. With $N$ calibration samples and sequence length $L$, the input activation is denoted as $\mathbf{X}$ with the shape of $(C_{\text{in}},N\times L)$. Layer-aware reconstruction error $\mathbf{E}$ is defined as the $\ell_{2}$ norm difference between output of dense and sparse layers:
$\mathbf{E}(\mathbf{\widehat{W}},\mathbf{X})=||\mathbf{WX}-\mathbf{\widehat{W}X}||_{2}^{2},$ (1)
where $\mathbf{\widehat{W}}$ is the element-wise product of $\mathbf{W}$ and a binary sparsity mask $\mathbf{M}(i,j)\in\{0,1\}$ of shape $(C_{\text{out}},C_{\text{in}})$, in context with mask selection and without weight reconstruction:
$\mathbf{E}(\mathbf{M},\mathbf{X})=||\mathbf{WX}-\mathbf{(W\odot M)X}||_{2}^{2}.$ (2)
The objective is to search for an optimal sparsity mask $\mathbf{M}$ where the pruning process reduces model complexity while preserving its predictive accuracy.
Taylor Expansion at Dense Networks. For a dense network $\mathbf{\widehat{W}_{\text{dense}}}$ at a local minimum, the reconstruction error can be expanded into its Taylor series with respect to $\mathbf{\widehat{W}}$, ignoring terms beyond second order:
$\mathbf{E}(\mathbf{\widehat{W}},\mathbf{X})=\mathbf{E}(\mathbf{\widehat{W}}_{\text{dense}},\mathbf{X})+\frac{\partial\mathbf{E}}{\partial\mathbf{W}}\Delta\mathbf{W}+\frac{1}{2}\Delta\mathbf{W}^{T}\mathbf{H}\Delta\mathbf{W}.$ (3)
Without a sparsity mask in dense model, we can simply assign all-one matrix to the mask M, thereby yielding the zeroth-order terms $\mathbf{E}(\mathbf{\widehat{W}_{\text{dense}}},\mathbf{X})=0$. The first-order derivative ${\partial\mathbf{E}}/\partial\mathbf{W}$ vanishes when training converged (Frantar and Alistarh 2023; Hassibi, Stork, and Wolff 1993), leaving only the computationally expensive second-order terms involving a large Hessian matrix which are challenging for layer-wise reconstruction and channel-wise independence assumption.
Block-Aware Rebuilder for Sparsity Mask
In this work, we depart from existing post-training pruning methods in three key aspects: Firstly, to address the layer-aware reconstruction problem that leads to exponentially accumulating biases, we adopt a block-aware reconstruction error and apply a divide-and-conquer strategy to mitigate errors and computational costs. Secondly, inspired by the pruning-and-growing111“Pruning” and “Growing” will both be used in the following text to refer to rebuild for sparsity masks. Pruning changes mask values from 1 to 0, whereas growing changes them from 0 to 1. operation (Mocanu et al. 2018; Zhang et al. 2024), we address the limitations of mask selection of pruning in dense networks due to the changeable significance of weights, by re-evaluating weight importance score in sparse networks and rebuilding the sparsity mask through targeted growth of salient weights and pruning of non-salient weights. Thirdly, our analysis reveals that with the advancement of LLMs, mask selection becomes increasingly critical in weight reconstruction. Thus, we prioritize the rebuilding of the sparsity mask and strip the reconstruction of weights. Base on these insights, we propose LLM-Barber, a Block-Aware Rebuilder for Sparsity Mask in One-Shot without any need for fine-tuning or retraining.
Block-Aware Reconstruction Error. Building on the definitions in Eq. (1) and Eq. (2), we define the block-aware reconstruction error for a Self-Attention or MLP block:
$\mathbf{E(M,X)}=||\mathbf{\text{Block}(W,X)}-\mathbf{\text{Block}(W\odot M,X)}||_{2}^{2}.$ (4)
Evaluating reconstruction error across blocks, denoted as $\mathbf{\text{Block}(\cdot)}$, allows us to achieve a globally optimal solution in Self-Attention and MLP blocks rather than layer-wise. This new block-aware reconstruction formulation offers a more effective and cohesive strategy for selecting sparsity masks.
Taylor Expansion at Sparse Networks. Migrating Eq. (3) at sparse networks $\mathbf{\widehat{W}_{\text{sparse}}}$ with an initialization sparsity mask $\mathbf{M_{i}}$, we can obtain the Taylor series expansion as:
$\mathbf{E(\widehat{W},X)}=\mathbf{E(\widehat{W}_{\text{sparse}},X)}+\mathbf{\frac{\partial E}{\partial W}}\Delta\mathbf{W}+\frac{1}{2}\Delta\mathbf{W}^{T}\mathbf{H}\Delta\mathbf{W}.$ (5)
The zeroth-order term of the Taylor expansion represents the reconstruction error after mask initialization:
$\mathbf{E(\widehat{W}_{\text{sparse}},X)}\!=||\mathbf{\text{Block}(W,X)}-\mathbf{\text{Block}(W\odot M_{i},X)}||_{2}^{2}.$ (6)
Assuming non-negligible zeroth-order terms, the first-order gradient in sparse networks remains significant even after convergence and can be efficiently accessed via PyTorch’s Autograd. First-order information provides computational efficiency and operates independently of any reconstruction error. Therefore, second-order terms can be omitted when significant gradients are present, leading to the following change in reconstruction error due to mask rebuilding:
$\Delta\mathbf{E}=(\partial\mathbf{E}/\partial\mathbf{W})\cdot\Delta\mathbf{W},$ (7)
which delineates the importance score of weights during sparsity mask rebuilding.
Pruning Metric. For a Self-Attention or MLP Block with weights $\mathbf{W}$, first-order information suffices for block-aware reconstruction error in sparse networks. The change in weight magnitude during sparsity mask adjustment matches the weight’s original magnitude ($|\Delta\mathbf{W}{ij}|=|\mathbf{W}{ij}|$). We thus assess the impact of mask rebuilding on reconstruction error by computing the product of the weight’s magnitude and its gradient. The importance score for $\mathbf{W}_{ij}$ is:
$\mathbf{S}_{ij}=|\mathbf{W}_{ij}|\cdot|(\partial\mathbf{E}/\partial\mathbf{W}_{ij})|,$ (8)
where ||$\cdot$|| ⋅ | represents the absolute value operator, and $\mathbf{E}$ is denoted as the block-aware reconstruction error:
$\mathbf{E(M,X)}=||\mathbf{\text{Block}(W,X)}-\mathbf{\text{Block}(W\odot M,X)}||_{2}^{2}.$ (9)
This method prioritizes weights with both substantial magnitudes and significant gradients, allowing for the preservation of critical weights while pruning less important ones.
Pruning Granularity. Choosing the right pruning granularity is crucial (Sun et al. 2023). Traditional methods operate on a layer-wise (Han et al. 2015), input-wise (Frantar and Alistarh 2023), or output-wise (Sun et al. 2023) basis, which mainly address the layer-aware reconstruction problem, known for its output channel independence. Wanda’s output-wise ranking yields superior results compared to other methods. However, in LLM-Barber’s block-aware framework, output channel independence is no longer applicable. Thus, LLM-Barber extends consideration to block-wise granularity, prioritizing all linear layers within a block. Our analysis of the four distinct granularity levels shows that optimal granularity depends on the specific sparse mask initialization, with detailed results discussed in the Ablation StudyAblation Study.
Mask Rebuilding. With the block-aware reconstruction error $\mathbf{E}$, gradient information, sparsity metric, and granularity established, we proceed to rebuild the sparsity mask for each layer. Consider a cluster of weights $\mathbf{W^{c}}$ under a specific sparsity granularity and its corresponding sparsity mask $\mathbf{M^{c}}$. We define the growing criterion and the pruning criterion:
$gi,gj=\text{argmax}\ |\mathbf{W^{c}}|\!\cdot\!|(\partial\mathbf{E}/\partial\mathbf{W^{c}})|,\text{if}\ \mathbf{M^{c}}_{gi,gj}=0,$ (10)
$pi,pj=\text{argmin}\ |\mathbf{W^{c}}|\!\cdot\!|(\partial\mathbf{E}/\partial\mathbf{W^{c}})|,\text{if}\ \mathbf{M^{c}}_{pi,pj}=1.$ (11)
The growing and pruning weights form a mask rebuilding pair, representing the interchange within the sparsity mask. The value of each pair is defined as the difference between the importance scores of the growing and pruning weights.
Our experiments show that LLM-Barber identifies varying proportions of salient weights depending on the sparse mask’s initialization method. To control the extent of mask rebuilding, we introduce a hyperparameter $\alpha$, called the mask rebuilding ratio. The number of mask rebuilding pairs $N$ is calculated as:
$N=\{\ i\ |\ \mathbf{S}^{\text{grow}}_{i}-\mathbf{S}^{\text{prune}}_{i}>0\}\cdot\alpha,$ (12)
where $\mathbf{S}^{\text{grow}}-\mathbf{S}^{\text{prune}}$ represents the value of the mask rebuilding pairs, where $\mathbf{S}^{\text{grow}}$ is arranged in descending order, and $\mathbf{S}^{\text{prune}}$ in ascending order. The subscript $i$ denotes the number of values that exceeds zero, indicating that the growing weight is more important than the pruning weight.
LLaMA1 LLaMA2 LLaMA3 OPT
Method 7B 13B 7B 13B 8B 6.7B 13B
Dense 5.677 5.091 5.472 4.884 6.136 10.86 10.13
Magnitude 17.26 20.14 16.03 6.827 205.5 9.7e2 1.2e4
w/ LLM-Barber 7.332 6.089 7.170 5.955 10.98 13.12 15.52
SparseGPT 7.201 6.194 7.005 6.036 9.399 11.59 11.15
SparseGPT w/o WR 7.545 6.311 7.413 6.134 9.994 13.13 15.76
w/ LLM-Barber 7.159 6.125 7.004 5.929 9.348 11.95 11.93
Wanda 7.254 6.152 6.920 5.972 9.821 11.98 11.93
w/ LLM-Barber 7.118 6.091 6.868 5.918 9.451 11.95 11.71
Table 1: WikiText-2 perplexity comparison for pruning LLM models at 50% sparsity rate. Bold results show improvements of integrating LLM-Barber. Underscored results indicate best performance in each LLM. WR represents weight reconstruction.
Input: Calibration samples $\mathbf{X}$, a block’s weights $\{\mathbf{W}^{l}\}^{L}_{l=1}$ and initial masks $\{\mathbf{M_{i}}^{l}\}^{L}_{l=1}$, mask rebuilding ratio $\alpha$.
Output: Rebuilt sparsity masks $\{\mathbf{M_{r}}^{l}\}^{L}_{l=1}$.
1:  $\mathbf{E_{i}}\leftarrow\text{Block}\mathbf{(W,X)}-\text{Block}\mathbf{(W\odot M_{i},X)}$
2:  $\{\mathbf{G}^{l}\}^{L}_{l=1}\leftarrow$ backpropagation for gradients via $\mathbf{E_{i}}$.
3:  for $l$ in $\{1,2,\dots,L\}$ do
4:     $\mathbf{M_{r}}^{l}\leftarrow\mathbf{M_{i}}^{l}$triangleright Initialize rebuilt sparsity mask.
5:     $\mathbf{S}^{l}\leftarrow|\mathbf{W}^{l}|\cdot|\mathbf{G}^{l}|$triangleright Obtain importance score.
[…]
7:     for $n$ in $\{1,2,\dots,N\}$ do
8:        Obtain growing index $gi,gj$ via Eq.(10)
9:        Obtain pruning index $pi,pj$ via Eq.(11)
10:        $\mathbf{M_{r}}^{l}_{pi,pj}=1$triangleright Weight Growing.
11:        $\mathbf{M_{r}}^{l}_{gi,gj}=0$triangleright Weight Pruning.
12:     end for
13:  end for
14:  $\mathbf{E_{r}}\leftarrow\text{Block}\mathbf{(W,X)}-\text{Block}\mathbf{(W\odot M_{r},X)}$
15:  Identify improvement by comparing error $\mathbf{E_{i}}$ and $\mathbf{E_{r}}$.
16:  return rebuilt sparsity masks $\{\mathbf{M_{r}}^{l}\}^{L}_{l=1}$.
Algorithm 1 Pseudocode of LLM-Barber.
Procedure
Like most post-training pruning methods (Han et al. 2015; Sun et al. 2023), LLM-Barber is executed within a single global LLM forward pass, with local backward passes for gradient computation in each block. Figure 2 illustrates the LLM-Barber workflow, which consists of four stages:
  • Sparsity Mask Initialization. We apply a post-training pruning technique to initialize a preliminary sparsity mask from the dense network.
  • Block-aware Reconstruction Error Computation. We use a block-aware reconstruction error to evaluate the discrepancy between the dense and sparse model outputs.
  • Back-propagation for Gradients. Gradients are automatically derived via back-propagation, and the product of weights and gradients serves as the pruning metric.
  • Sparsity Mask Rebuilding. Masks are sorted based on the pruning metric, unpruned weights in ascending order while pruned weights in descending order. We rebuild the weight masks by growing newly significant weights and pruning those became non-salient with a certain percentage named mask rebuilding ratio.
Ultimately, we recalculated the block-aware reconstruction error to assess the enhancement brought by our LLM-Barber. Experimental results reveal a substantial decrease in error, markedly decelerating the rate of error accumulation.
Structured N:M Sparsity
While LLM-Barber primarily targets unstructured sparsity, it can be adapted for structured N:M sparsity. Here, groups of M weights are pruned to retain only N non-zero weights. During mask rebuilding, LLM-Barber divides each M-group into N pairs (each pair contains one pruned and one non-pruned weight), then sorts these pairs by output channel to determine which are mask rebuilding pairs. This method creates an optimal sparse mask, leveraging N:M sparsity while maintaining model performance.
Experiment
Experiment Settings
Setup. LLM-Barber is implemented in Pytorch and utilized public model checkpoints from the HuggingFace library 222huggingface.co/meta-llama, huggingface.co/huggyllama on a single 80GB NVIDIA A100 GPU. After mask initilization, LLM-Barber uniformly rebuilds sparsity masks in sequence, performing in one-shot without any fine-tuning.
Models & Datasets. LLM-Barber is evaluated on the LLaMA family, including LLaMA-7B/13B (Touvron et al. 2023a), LLaMA2-7B/13B (Touvron et al. 2023b), and LLaMA3-8B (AI@Meta 2024), as well as the OPT model series: OPT-6.7B/13B (Zhang et al. 2022). Notably, LLM-Barber is broadly applicable to any Transformer-based LLMs with Self-Attention and MLP blocks. Following previous works, we use 128 segments of 2048 tokens from the C4 dataset (Raffel et al. 2020) for mask rebuilding.
Evaluation. To comprehensively assess LLM-Barber, we conduct rigorous evaluations on perplexity and zero-shot accuracy. Perplexity is measured on the validation sets of benchmarks such as WikiText-2 (Merity et al. 2017), PTB (Marcus et al. 1994), and C4 (Raffel et al. 2020). Zero-shot accuracy is assessed using the EleutherAI LM Harness (Gao et al. 2023) across six benchmarks: BoolQ (Clark et al. 2019), RTE (Wang et al. 2018), HellaSwag (Zellers et al. 2019), ARC Easy and Challenge (Clark et al. 2018), and OpenbookQA (Mihaylov et al. 2018).
Baselines. The results of LLM-Barber are compared with the following established post-training pruning methods:
  • Magnitude Pruning (Han et al. 2015) eliminates weights based only on their magnitudes;
  • SparseGPT (Frantar and Alistarh 2023) identifies weights importance by using second-order information;
  • Wanda (Sun et al. 2023) determines weights to be pruned by the weight magnitude multiplied by input activation.
Main Results
Unstructured Sparsity. LLM-Barber effectively prunes the LLaMA and OPT models, achieving 50% unstructured sparsity without requiring supplementary weight reconstruction, as detailed in Table 1. LLM-Barber demonstrates the capability to rebuild the sparsity masks initialized by other pruning methods in a single forward pass, significantly outperforming conventional pruning baselines. In the LLaMA3-8B model, LLM-Barber creates a new sparsity mask that reduces perplexity to 9.451, a substantial improvement over the Wanda baseline of 9.821. Notably, LLM-Barber achieves robust improvements even with poorly performing initial sparsity masks, such as magnitude pruning, where it reduces perplexity from 205.5 to 10.98, which is an impressive and substantial enhancement.
Sparsity 60% 70% 80% 90%
Magnitude 3.39e4 1.62e6 8.55e7 2.26e7
w/ LLM-Barber 28.14 2.08e2 1.09e3 7.55e4
Wanda 23.57 1.28e2 8.54e2 1.28e4
w/ LLM-Barber 22.04 1.07e2 5.70e2 7.07e3
Table 2: WikiText-2 perplexity performance for pruning LLaMA3-8B at varying sparsity rate.
Varying Sparsity Levels. We conduct experiments on varying sparsity levels for unstructured pruning in LLaMA3-8B as shown in Table 2. LLM-Barber consistently increases perplexity across all initialization methods, with magnitude pruning showing the most significant improvement.
Method Sparsity V1-7B V1-13B V2-7B
Magnitude 4:8 16.83 13.72 15.91
w/ LLM-Barber 4:8 8.852 7.137 9.345
SparseGPT 4:8 8.608 7.437 8.495
w/ LLM-Barber 4:8 8.191 7.085 8.003
Magnitude 2:4 42.56 18.32 37.76
w/ LLM-Barber 2:4 11.04 9.006 13.47
SparseGPT 2:4 11.55 9.116 10.94
w/ LLM-Barber 2:4 10.14 8.517 9.806
Table 3: WikiText-2 perplexity comparison for pruning LLaMA family with structured N:M pattern.
Structured N:M Sparsity. In contrast to unstructured sparsity, employing N:M fine-grained sparsity can provide more tangible acceleration benefits when leveraging NVIDIA Ampere’s sparse tensor cores (Choquette et al. 2021).
Model Method BoolQ RTE HellaSwag ARC-e ARC-c OBQA Mean
LLaMA-7B Dense 75.08 66.79 56.96 75.29 41.89 34.40 58.40
Magnitude 54.61 54.51 45.47 58.75 33.45 22.60 44.89
w/ LLM-Barber 71.47 59.93 50.63 69.11 35.41 28.20 52.46
Wanda 70.98 55.23 51.90 69.44 36.86 28.60 52.17
w/LLM-Barber 73.12 56.68 51.80 70.32 36.95 28.80 52.95
LLaMA2-13B Dense 80.58 65.34 60.05 79.38 48.46 35.20 61.50
Magnitude 70.53 55.96 54.42 57.68 38.40 27.80 50.79
w/ LLM-Barber 81.10 60.29 55.67 75.34 40.28 31.20 57.31
Wanda 80.95 60.28 56.98 76.30 42.26 31.20 57.99
w/ LLM-Barber 80.98 62.82 55.99 76.39 42.13 31.40 58.29
LLaMA3-8B Dense 81.35 69.68 60.19 80.09 50.60 34.80 62.79
Magnitude 42.87 53.07 29.85 46.59 25.09 22.00 36.57
w/ LLM-Barber 72.72 54.87 51.00 72.05 37.46 27.40 52.58
Wanda 78.41 60.29 51.20 71.38 40.10 29.40 55.13
w/LLM-Barber 78.59 61.37 51.01 71.46 39.85 29.00 55.21
Table 4: Zero-shot performance comparison for pruning LLaMA series on six evaluated tasks at 50% sparsity rate. Bold results show improvements of integrating LLM-Barber. Underscored results indicate the best performance in each zero-shot tasks.
Therefore, we also evaluate the effectiveness of our LLM-Barber in partial LLaMA models on the N:M fine-grained sparsity pattern as shown in Table 3.
Zero-shot Performance. Following previous works (Frantar and Alistarh 2023; Sun et al. 2023), we evaluated the LLaMA models on six diverse zero-shot tasks. The results are summarized in Table 4, where models are pruned to unstructured 50% sparsity. Averaging the accuracy across the six evaluated tasks, it becomes apparent that LLM-Barber possesses the capability to identify a more effective network than those obtained via the initialization methods. For tasks such as BoolQ, RTE, and ARC-e, LLM-Barber consistently outperforms the baseline pruning techniques across the entire LLaMA model suite. However, it is worth noting that there is no single universally superior performer for the remaining tasks in the evaluation set, with the initial pruning methods sometimes matching or even marginally exceeding the results obtained through LLM-Barber.
Mask Rebuilding Ratio Selection
Figure 3: Relationship between importance score distribution of mask rebuilding pair and WikiText-2 perplexity results at varying pruning granularities in LLaMA-7B. The green dashed line indicates the best choice of mask rebuilding ratio.
A critical aspect of the LLM-Barber method lies in determining the optimal mask rebuilding ratio to achieve peak accuracy. One effective strategy involves analyzing the distribution of value magnitudes within the mask rebuilding pairs, corresponding to differences between the growing and pruning importance score. Fortunately, this distribution often exhibits a distinct pattern of outliers, facilitating rapid identification of an appropriate mask rebuilding ratio.
We have plotted the score distribution of mask rebuilding pairs for various pruning granularity and initialization methods on LLaMA-7B, as well as perplexity performance corresponding to mask rebuilding ratios in Figure 3. The results reveal a strong correlation between the distribution of outliers in mask rebuilding pairs and the optimal mask rebuilding ratio. For instance, outliers are significantly distributed within the top 10% with Magnitude pruning, thus selecting a 10% mask rebuilding ratio yields the overall optimal solution. Similarly, with the initialization method Wanda, a significant outlier distribution around the 1% mark corresponds to optimal results near a 1% mask rebuilding ratio. Thus, LLM-Barber can preemptively narrow the search range for the optimal mask rebuilding ratio by analyzing outlier distributions. This approach allows LLM-Barber to adapt flexibly to various reconstruction masks without extensive searching.
It is worth noting that LLM-Barber identifies a larger proportion of outliers in less effective initial masks, which corresponds to a more substantial rebuilding of the mask. This is why LLM-Barber provides a more significant improvement for masks with poorer initial sparsity mask with a more aggressive mask rebuilding strategy.
Ablation Study
Given the outstanding potential of LLM-Barber, we have comprehensively analyzed the impact of three critical factors to assess the robustness and effectiveness of our LLM-Barber method: pruning granularity, pruning metric, and calibration data size. These factors were chosen to gain a deeper understanding of how different configurations influence the performance of pruned models and to demonstrate LLM-Barber’s versatility across various settings.
Pruning Granularity. LLM-Barber dynamically selects different pruning granularities to adapt to varying initialization methods. In this paper, we evaluated the impact of four levels of granularities: block-wise, layer-wise, input-wise, and output-wise pruning, as shown in Table 5. For magnitude pruning, the block-wise granularity yields the best performance, while the output-wise granularity delivers the lowest perplexity for Wanda pruning. By tailoring the pruning granularity to the particular pruning approach, LLM-Barber can consistently achieve optimal model performance.
Method Pruning Granularity
w/ LLM-Barber Block Layer Input Output
Magnitude (205.5) 10.98 11.06 72.48 11.81
SparseGPT (9.994) 9.380 9.348 9.418 9.567
Wanda (9.821) 9.626 9.633 9.849 9.451
Table 5: Ablation of pruning granularity in LLaMA3-8B. Perplexity in brackets show the WikiText-2 score without LLM-Barber. Bold results show best granularity of this row.
Pruning Metric. We analyze the effect of different pruning metrics, focusing on weight magnitude, gradient magnitude, and the product of weight and gradient. As seen in Table 6, the product of weight and gradient consistently outperforms other metrics, achieving the lowest perplexity of 10.98 for Magnitude, 9.348 for SparseGPT, and 9.451 for Wanda. This confirms the effectiveness of our product-based pruning metric across various methods.
Method Pruning Metric
w/ LLM-Barber $|\mathbf{W}|$ $|\mathbf{\partial E/\partial W}|$ $|\mathbf{W}|\cdot|\mathbf{\partial E/\partial W}|$
Magnitude (205.5) 186.5 14.43 10.98
SparseGPT (9.994) 9.544 9.457 9.348
Wanda (9.821) 9.880 9.699 9.451
Table 6: Ablation of pruning metric in LLaMA3-8B. Bold results support the pruning metric $|\mathbf{W}|\cdot|\mathbf{\partial E/\partial W}|$.
Calibration Data Size. We explore how varying the size of calibration data influences the performance of LLM-Barber. Figure 4 demonstrates that as the calibration sample size increases, LLM-Barber maintains robust performance. Notably, with more than four samples, our method achieves better perplexity than SparseGPT and Wanda. LLM-Barber outperforms SparseGPT even with just a single sample, underscoring its robustness across different sample sizes.
Figure 4: Ablation of calibration data size in LLaMA3-8B. LLM-Barber is robust across varying calibration data size.
Conclusion
We propose LLM-Barber (Block-Aware Rebuilder for Sparsity Mask in One-Shot), a novel framework that rebuilds sparsity mask to optimize LLM post-training pruning. By integrating a block-aware approach across Self-Attention and MLP blocks, LLM-Barber effectively reallocates sparsity to improve accuracy without the need for extensive fine-tuning. Specifically, LLM-Barber identifies novel importance score after mask initialization and rebuilds the sparsity mask with mask rebuilding pairs, simultaneously applying new sparsity masks to weights that have become less critical, thereby optimizing overall model performance. By utilizing the novel pruning metric as the product of weights and gradients, our approach enables precise and efficient reallocation of sparsity mask. Extensive experiments on pruning LLaMA series models demonstrate that LLM-Barber achieves state-of-the-art results in both perplexity and zero-shot performance within the domain of post-training pruning.

### v2

@@
-LLM-Barber: Block-Aware Rebuilder for Sparsity Mask in One-Shot for Large Language Models
+Graziano Chesi2, Ngai Wong2, Hao Yu1
 Abstract
-Large language models (LLMs) have grown significantly in scale, leading to a critical need for efficient model pruning techniques.
-Existing post-training pruning techniques primarily focus on measuring weight importance on converged dense models to determine salient weights to retain.
-However, they often overlook the changes in weight importance during the pruning process, which can lead to performance degradation in the pruned models.
+Large language models (LLMs) have seen substantial growth, necessitating efficient model pruning techniques.
+Existing post-training pruning methods primarily measure weight importance in converged dense models, often overlooking changes in weight significance during the pruning process, leading to performance degradation.
 To address this issue, we present LLM-Barber (Block-Aware Rebuilder for Sparsity Mask in One-Shot), a novel one-shot pruning framework that rebuilds the sparsity mask of pruned models without any retraining or weight reconstruction.
-LLM-Barber incorporates block-aware error optimization across Self-Attention and MLP blocks, ensuring global performance optimization.
-Inspired by the recent discovery of prominent outliers in LLMs, LLM-Barber introduces an innovative pruning metric that identifies weight importance using weights multiplied by gradients.
-Our experiments show that LLM-Barber can efficiently prune models like LLaMA and OPT families with 7B to 13B parameters on a single A100 GPU in just 30 minutes, achieving state-of-the-art results in both perplexity and zero-shot performance across various language benchmarks.
-Code is available at https://github.com/YupengSu/LLM-Barber.
-Introduction
-Large language models (LLMs) have become a cornerstone in natural language processing (NLP) due to their impressive performance on various tasks.
-However, as these models increase in size and complexity, their deployment poses significant challenges due to extensive computational and storage demands.
-For instance, models such as GPT-175B (Brown et al.
-2020), with 175 billion parameters, require vast resources, making it impractical for many applications.
-Therefore, efficient model compression strategies are crucial for deploying these powerful models in practical applications.
-Model compression techniques commonly employ quantization and pruning to enhance model efficiency.
-Quantization reduces the precision of model parameters, while pruning removes less critical parameters.
-Traditional pruning methods, such as magnitude-based pruning (Han et al.
-2015), directly trim weights based on their absolute values.
-While effective for smaller models, these methods often struggle with large-scale LLMs, resulting in suboptimal sparsity due to their inability to capture the complex interactions and importance of weights in such massive scalability.
-To address these limitations, recent post-training pruning techniques such as SparseGPT and Wanda have emerged.
+LLM-Barber incorporates block-aware error optimization across Self-Attention and MLP blocks, facilitating global performance optimization.
+We are the first to employ the product of weights and gradients as a pruning metric in the context of LLM post-training pruning.
+This enables accurate identification of weight importance in massive models and significantly reduces computational complexity compared to methods using second-order information.
+Our experiments show that LLM-Barber efficiently prunes models from LLaMA and OPT families (7B to 13B) on a single A100 GPU in just 30 minutes, achieving state-of-the-art results in both perplexity and zero-shot performance across various language benchmarks.
+Index Terms:
+Large Language Models, Post-training Pruning, Sparsity Mask Rebuilding, Low-resource Deployment
+I Introduction
+LLMs have become foundational in artificial intelligence due to their impressive performance on various tasks.
+However, the increasing size and complexity of models, such as GPT-175B [1] with 175 billion parameters, pose significant challenges related to extensive computational and storage demands.
+Consequently, efficient model compression strategies are crucial for enabling the practical deployment of these powerful models.
 Current pruning methods face two major challenges.
-First, as depicted in Figure 1(a), traditional layer-aware pruning methods focus on individual layers and neglect inter-layer dependencies within, leading to higher error accumulation (represented by blue arrows).
-In contrast, block-aware pruning, by considering groups of layers, captures inter-layer interactions to reduce error accumulation (represented by orange arrows).
-Second, as shown in Figure 1(b), conventional methods typically build the pruning mask once, ignoring the changes of weight significance in post-pruning stage.
-This oversight can lead to improper identification of salient weights, resulting in performance degradation.
+First, traditional layer-aware pruning methods focus on individual layers, neglecting inter-layer dependencies, which increases error accumulation (blue arrows in Figure 1(a)).
+In contrast, block-aware pruning considers groups of layers, effectively reducing this error (orange arrows).
+Second, as shown in Figure 1(b), conventional methods typically build the pruning mask only once, ignoring the changes of weight significance in post-pruning stage.
+This oversight can lead to improper identification of salient weights and subsequent performance degradation.
 Figure 1: The benefits of integrating LLM-Barber into the pruning process: (a) Transition from the layer-aware to block-aware error accumulation to achieve an optimized global solution.
 (b) Rebuilding sparsity mask using a novel pruning metric based on weights multiplied by gradients.
-To address these limitations, we propose LLM-Barber, a novel and straightforward approach designed to rebuild sparsity mask of pruned networks without requiring for retraining or weight reconstruction.
-Firstly, unlike layer-aware methods that are confined to local optimization and thus prone to significant error accumulation, as depicted by the blue arrows in Figure 1(a), LLM-Barber integrates pruning across both Self-Attention and MLP block.
-This approach mitigates error accumulation, as evidenced by the lighter orange arrows, facilitating global optimization and improved model performance.
+To address these limitations, we propose LLM-Barber, a novel and straightforward approach rebuild sparsity mask of pruned networks without requiring for retraining or weight reconstruction.
+Figure 2: The workflow of LLM-Barber.
+(a) illustrates the process of block-aware reconstruction error and gradient calculation for each linear weight.
+(b) shows pruning metric computation and sparsity mask rebuilding.
+Firstly, unlike layer-aware methods that prioritize local optimization and are prone to error accumulation (illustrated by the blue arrows in Figure 1(a)), LLM-Barber integrates pruning across Self-Attention and MLP blocks.
+This approach mitigates error accumulation, as evidenced by the lighter orange arrows, promoting more global optimization.
 Secondly, LLM-Barber identifies weights that, although initially non-salient without a sparsity mask, gain significance in post-pruning.
 As shown in Figure 1(b), varying color shades represent the relative importance scores of different weights.
-LLM-Barber accurately identifies and rebuilds masks for high-score (deeply shaded) weights while pruning newly identified low-score (lightly shaded) weights.
-Thirdly, LLM-Barber utilizes a first-order Taylor series as pruning metric, which leverages weight-gradient multiplication for precise rebuilding of the sparsity mask, thereby enabling accurate pruning decisions and significantly reducing computational complexity compared to methods using second-order information.
-Finally, our one-shot approach surpasses iterative fine-tuning in efficiency while maintaining comparable accuracy.
-By rebuilding sparsity masks of pruned model at once, LLM-Barber offers a faster and more precise pruning solution.
-We validate the effectiveness of LLM-Barber through extensive experiments on the LLaMA and OPT families.
-Our method consistently outperforms existing post-training pruning techniques in both perplexity and zero-shot performance, establishing LLM-Barber as a state-of-the-art approach in this domain.
+LLM-Barber accurately rebuilds masks for high-score weights (deeply shaded) while pruning newly identified low-score weights (lightly shaded).
+thirdly, LLM-Barber employs the product of weights and gradients as pruning metric.
+While this metric has been applied in other contexts [2, 3], we are the first to leverage it in the context of post-training compression for LLMs.
+This enables precise rebuilding of the sparsity mask and significantly reduces computational complexity compared to methods [4] using second-order information.
+Finally, our one-shot approach efficiently rebuilds sparsity masks for pruned models, achieving comparable accuracy to iterative fine-tuning.
+LLM-Barber demonstrates its effectiveness through experiments on the LLaMA and OPT families, consistently outperforms existing post-training techniques in both perplexity and zero-shot tasks.
 To sum up, the key contributions are fourfold:
@@
 This targeted rebuilding enhances overall model performance by optimally reallocating sparsity.
-  • Innovative Pruning Metric: We propose an innovative pruning metric based on the product of weights and gradients, leveraging first-order Taylor series for importance evaluation to reduce computational complexity compared to second-order Hessian-based approaches.
+  • Innovative Pruning Metric: We are the first to apply the product of weights and gradients as a pruning metric within the context of LLM post-training pruning, leveraging first-order Taylor series for importance evaluation to reduce computational complexity compared to second-order Hessian-based approaches.
   • Versatility and Efficiency: LLM-Barber demonstrates its efficiency and effectiveness across various pruning techniques, consistently achieving state-of-the-art performance in perplexity and zero-shot tasks, thereby establishing new yardsticks in LLM post-training pruning.
-Related Work
-LLMs Pruning and Sparsity
-Network pruning reduces deep neural networks by removing unnecessary weights.
-For LLMs, pruning methods are divided into parameter efficient fine-tuning and post-training approaches.
-Parameter Efficient Fine-tuning (PEFT) begins with a initialized sparse network and refines it through iterative processes (Liu et al.
-2019).
-LoRA (Hu et al.
-2021) adapts pre-trained models to specific tasks or domains by injecting trainable rank decomposition matrices.
-Dynamic Spase No Training (Zhang et al.
-2024) minimizes the reconstruction error by iteratively pruning and growing weights.
-However, fine-tuning requires ample data and often leads to performance decline.
-Recent studies advance toward one-shot post-training pruning, showing substantial improvements.
-Post-training pruning removes weights from a pre-trained model.
-SparseGPT (Frantar and Alistarh 2023) uses Hessian-based metrics and subsequent residual weight updates, while Wanda (Sun et al.
-2023) introduces a first-order pruning metric using weight-activation products.
-LLM-Barber employs a block-aware reconstruction approach and rebuilds masks with a novel pruning metric.
-Model Compression Strategy
+II Related Work
+II-A LLMs Pruning and Sparsity
+Network pruning reduces deep neural networks by eliminating unnecessary weights, with methods categorized into parameter-efficient fine-tuning (PEFT) and post-training approaches.
+PEFT begins with an initialized sparse network and refines it through iterative processes [5].
+LoRA [6] adapts pre-trained models to specific tasks by injecting trainable rank decomposition matrices.
+Dynamic Spase No Training [7] prunes and grows weights to minimize reconstruction error.
+However, fine-tuning often requires ample data and can degrade performance.
+Post-training, on the other hand, removes weights from a pre-trained model.
+SparseGPT [4] uses Hessian-based metrics and subsequent residual weight updates, Wanda [8] introduces a first-order metric using weight-activation products.
+LLM-Barber employs a block-aware reconstruction strategy and rebuilds masks with a novel pruning metric.
+II-B Model Compression Strategy
 Compression is key to reducing the memory and computational demands of model.
-This work highlights the block-aware compression strategy over traditional layer-aware approaches.
-Layer-aware compression strategy began with Optimal Brain Damage (LeCun, Denker, and Solla 1989) and Optimal Brain Surgeon (Hassibi, Stork, and Wolff 1993).
-Building on these foundations, recent works like GPTQ (Frantar et al.
-2023) further enhance layer-aware quantization using second-order information.
-Block-aware compression strategy generally offer better accuracy recovery in pruned models compared to layer-aware methods.
-For instance, APTQ (Guan et al.
-2024) applies global quantization to attention mechanisms, enhancing model robustness, while BESA (Xu et al.
-2024) uses block-wise sparsity allocation.
-Our method leverages block-aware pruning to optimize global performance across Self-Attention and MLP blocks, effectively balancing efficiency and accuracy.
-Block-Aware Rebuilder for Sparsity Mask
-Figure 2: The workflow of LLM-Barber.
-(a) illustrates the process of block-aware reconstruction error calculation and gradient back-propagation for each linear weight.
-(b) shows pruning metric computation and procedure for rebuilding sparsity mask.
-Preliminaries
-LLM pruning removes weights from dense networks to minimize output discrepancies, which is computationally intensive cross large-scale models, leading to address a layer-aware reconstruction problem (Hassibi, Stork, and Wolff 1993).
+Layer-aware strategies, originating from Optimal Brain Damage [9] and Optimal Brain Surgeon [10], have been enhanced by recent works like GPTQ [11] using second-order information.
+However, block-aware strategies typically yield better accuracy recovery.
+For instance, APTQ [12] applies global quantization to attention mechanisms, BESA [13] uses block-wise sparsity allocation.
+Our method leverages block-aware pruning to optimize global performance across blocks, effectively balancing efficiency and accuracy.
+III LLM-Barber
+III-A Preliminaries
+LLM pruning removes weights from dense networks to minimize output discrepancies, which is computationally intensive cross large-scale models, leading to address a layer-aware reconstruction problem [10].
 This section reviews and reanalyses layer-aware reconstruction error and Taylor expansion at dense networks.
@@
 Layer-aware reconstruction error $\mathbf{E}$ is defined as the $\ell_{2}$ norm difference between output of dense and sparse layers:
-$\mathbf{E}(\mathbf{\widehat{W}},\mathbf{X})=||\mathbf{WX}-\mathbf{\widehat{W}X}||_{2}^{2},$ (1)
+$\mathbf{E}=||\mathbf{WX}-\mathbf{\widehat{W}X}||_{2}^{2},$ (1)
 where $\mathbf{\widehat{W}}$ is the element-wise product of $\mathbf{W}$ and a binary sparsity mask $\mathbf{M}(i,j)\in\{0,1\}$ of shape $(C_{\text{out}},C_{\text{in}})$, in context with mask selection and without weight reconstruction:
-$\mathbf{E}(\mathbf{M},\mathbf{X})=||\mathbf{WX}-\mathbf{(W\odot M)X}||_{2}^{2}.$ (2)
-The objective is to search for an optimal sparsity mask $\mathbf{M}$ where the pruning process reduces model complexity while preserving its predictive accuracy.
+$\mathbf{E}=||\mathbf{WX}-\mathbf{(W\odot M)X}||_{2}^{2}.$ (2)
+The goal is to find an optimal sparsity mask $\mathbf{M}$ that reduces model complexity while preserving predictive accuracy.
 Taylor Expansion at Dense Networks.
 For a dense network $\mathbf{\widehat{W}_{\text{dense}}}$ at a local minimum, the reconstruction error can be expanded into its Taylor series with respect to $\mathbf{\widehat{W}}$, ignoring terms beyond second order:
-$\mathbf{E}(\mathbf{\widehat{W}},\mathbf{X})=\mathbf{E}(\mathbf{\widehat{W}}_{\text{dense}},\mathbf{X})+\frac{\partial\mathbf{E}}{\partial\mathbf{W}}\Delta\mathbf{W}+\frac{1}{2}\Delta\mathbf{W}^{T}\mathbf{H}\Delta\mathbf{W}.$ (3)
-Without a sparsity mask in dense model, we can simply assign all-one matrix to the mask M, thereby yielding the zeroth-order terms $\mathbf{E}(\mathbf{\widehat{W}_{\text{dense}}},\mathbf{X})=0$.
-The first-order derivative ${\partial\mathbf{E}}/\partial\mathbf{W}$ vanishes when training converged (Frantar and Alistarh 2023; Hassibi, Stork, and Wolff 1993), leaving only the computationally expensive second-order terms involving a large Hessian matrix which are challenging for layer-wise reconstruction and channel-wise independence assumption.
-Block-Aware Rebuilder for Sparsity Mask
+$\mathbf{E}=\mathbf{E_{d}}+\frac{\partial\mathbf{E}}{\partial\mathbf{W}}\Delta\mathbf{W}+\frac{1}{2}\Delta\mathbf{W}^{T}\mathbf{H}\Delta\mathbf{W}.$ (3)
+Without a sparsity mask in dense model, we can simply assign all-one matrix to the mask M, thereby yielding the zeroth-order terms $\mathbf{E_{d}}=0$.
+The first-order derivative ${\partial\mathbf{E}}/\partial\mathbf{W}$ vanishes when training converged [4, 10], leaving only the computationally expensive second-order terms involving a large Hessian matrix which are challenging for layer-wise reconstruction and channel-wise independence assumption.
+III-B Block-Aware Rebuilder for Sparsity Mask
 In this work, we depart from existing post-training pruning methods in three key aspects: Firstly, to address the layer-aware reconstruction problem that leads to exponentially accumulating biases, we adopt a block-aware reconstruction error and apply a divide-and-conquer strategy to mitigate errors and computational costs.
-Secondly, inspired by the pruning-and-growing111“Pruning” and “Growing” will both be used in the following text to refer to rebuild for sparsity masks.
-Pruning changes mask values from 1 to 0, whereas growing changes them from 0 to 1.
-operation (Mocanu et al.
-2018; Zhang et al.
-2024), we address the limitations of mask selection of pruning in dense networks due to the changeable significance of weights, by re-evaluating weight importance score in sparse networks and rebuilding the sparsity mask through targeted growth of salient weights and pruning of non-salient weights.
+Secondly, inspired by the pruning-and-growing111“Pruning” and “Growing” will both be used in following text to refer to rebuild sparsity masks.
+Pruning changes masks from 1 to 0, whereas growing changes them from 0 to 1.
+operation [14, 7], we address the limitations of mask selection of pruning in dense networks due to the changeable significance of weights, by re-evaluating weight importance score in sparse networks and rebuilding the sparsity mask through targeted growth of salient weights and pruning of non-salient weights.
 Thirdly, our analysis reveals that with the advancement of LLMs, mask selection becomes increasingly critical in weight reconstruction.
@@
 Block-Aware Reconstruction Error. Building on the definitions in Eq. (1) and Eq. (2), we define the block-aware reconstruction error for a Self-Attention or MLP block:
-$\mathbf{E(M,X)}=||\mathbf{\text{Block}(W,X)}-\mathbf{\text{Block}(W\odot M,X)}||_{2}^{2}.$ (4)
-Evaluating reconstruction error across blocks, denoted as $\mathbf{\text{Block}(\cdot)}$, allows us to achieve a globally optimal solution in Self-Attention and MLP blocks rather than layer-wise.
-This new block-aware reconstruction formulation offers a more effective and cohesive strategy for selecting sparsity masks.
+$\mathbf{E}=||\mathbf{\text{Block}(W,X)}-\mathbf{\text{Block}(W\odot M,X)}||_{2}^{2}.$ (4)
+[…]
 Taylor Expansion at Sparse Networks.
@@
 (3) at sparse networks $\mathbf{\widehat{W}_{\text{sparse}}}$ with an initialization sparsity mask $\mathbf{M_{i}}$, we can obtain the Taylor series expansion as:
-$\mathbf{E(\widehat{W},X)}=\mathbf{E(\widehat{W}_{\text{sparse}},X)}+\mathbf{\frac{\partial E}{\partial W}}\Delta\mathbf{W}+\frac{1}{2}\Delta\mathbf{W}^{T}\mathbf{H}\Delta\mathbf{W}.$ (5)
-The zeroth-order term of the Taylor expansion represents the reconstruction error after mask initialization:
-$\mathbf{E(\widehat{W}_{\text{sparse}},X)}\!=||\mathbf{\text{Block}(W,X)}-\mathbf{\text{Block}(W\odot M_{i},X)}||_{2}^{2}.$ (6)
+$\mathbf{E}=\mathbf{E_{s}}+\mathbf{\frac{\partial E}{\partial W}}\Delta\mathbf{W}+\frac{1}{2}\Delta\mathbf{W}^{T}\mathbf{H}\Delta\mathbf{W}.$ (5)
+The zeroth-order term $\mathbf{E_{s}}$ of the Taylor expansion at sparse networks represents the reconstruction error after mask initialization:
+$\mathbf{E_{s}}=||\mathbf{\text{Block}(W,X)}-\mathbf{\text{Block}(W\odot M_{i},X)}||_{2}^{2}.$ (6)
 Assuming non-negligible zeroth-order terms, the first-order gradient in sparse networks remains significant even after convergence and can be efficiently accessed via PyTorch’s Autograd.
@@
 $\mathbf{S}_{ij}=|\mathbf{W}_{ij}|\cdot|(\partial\mathbf{E}/\partial\mathbf{W}_{ij})|,$ (8)
-where ||$\cdot$|| ⋅ | represents the absolute value operator, and $\mathbf{E}$ is denoted as the block-aware reconstruction error:
-$\mathbf{E(M,X)}=||\mathbf{\text{Block}(W,X)}-\mathbf{\text{Block}(W\odot M,X)}||_{2}^{2}.$ (9)
+where ||$\cdot$|| ⋅ | represents the absolute value operator, and $\mathbf{E}$ denotes the block-aware reconstruction error:
+$\mathbf{E}=||\mathbf{\text{Block}(W,X)}-\mathbf{\text{Block}(W\odot M,X)}||_{2}^{2}.$ (9)
 This method prioritizes weights with both substantial magnitudes and significant gradients, allowing for the preservation of critical weights while pruning less important ones.
 Pruning Granularity.
-Choosing the right pruning granularity is crucial (Sun et al.
-2023).
-Traditional methods operate on a layer-wise (Han et al.
-2015), input-wise (Frantar and Alistarh 2023), or output-wise (Sun et al.
-2023) basis, which mainly address the layer-aware reconstruction problem, known for its output channel independence.
+Choosing the right pruning granularity is crucial [8].
+Traditional methods operate on a layer-wise [15], input-wise [4], or output-wise [8] basis, which mainly address the layer-aware reconstruction problem, known for its output channel independence.
 Wanda’s output-wise ranking yields superior results compared to other methods.
@@
 Thus, LLM-Barber extends consideration to block-wise granularity, prioritizing all linear layers within a block.
-Our analysis of the four distinct granularity levels shows that optimal granularity depends on the specific sparse mask initialization, with detailed results discussed in the Ablation StudyAblation Study.
+Our analysis of the four distinct granularity levels shows that optimal granularity depends on the specific sparse mask initialization, with detailed results discussed in the Ablation Study IV-D.
 Mask Rebuilding.
-With the block-aware reconstruction error $\mathbf{E}$, gradient information, sparsity metric, and granularity established, we proceed to rebuild the sparsity mask for each layer.
-Consider a cluster of weights $\mathbf{W^{c}}$ under a specific sparsity granularity and its corresponding sparsity mask $\mathbf{M^{c}}$.
+With the block-aware reconstruction error, gradient information, sparsity matrix, and granularity established, we proceed to rebuild the sparsity mask for each layer.
+Consider a cluster of weights $\mathbf{W^{c}}$ under a specific sparsity granularity and its corresponding sparsity mask $\mathbf{M^{c}}$ and pruning metric $\mathbf{S^{c}}$.
 We define the growing criterion and the pruning criterion:
-$gi,gj=\text{argmax}\ |\mathbf{W^{c}}|\!\cdot\!|(\partial\mathbf{E}/\partial\mathbf{W^{c}})|,\text{if}\ \mathbf{M^{c}}_{gi,gj}=0,$ (10)
-$pi,pj=\text{argmin}\ |\mathbf{W^{c}}|\!\cdot\!|(\partial\mathbf{E}/\partial\mathbf{W^{c}})|,\text{if}\ \mathbf{M^{c}}_{pi,pj}=1.$ (11)
+$gi,gj=\text{argmax}\ \mathbf{S^{c}},\ \text{if}\ \mathbf{M^{c}}_{gi,gj}=0,$ (10)
+$pi,pj=\text{argmin}\ \mathbf{S^{c}},\ \text{if}\ \mathbf{M^{c}}_{pi,pj}=1.$ (11)
 The growing and pruning weights form a mask rebuilding pair, representing the interchange within the sparsity mask.
 The value of each pair is defined as the difference between the importance scores of the growing and pruning weights.
-Our experiments show that LLM-Barber identifies varying proportions of salient weights depending on the sparse mask’s initialization method.
-To control the extent of mask rebuilding, we introduce a hyperparameter $\alpha$, called the mask rebuilding ratio.
+Our experiments show that LLM-Barber identifies varying proportions of salient weights based on the sparse mask’s initialization method.
+To regulate mask rebuilding, we introduce a hyperparameter $\alpha$, known as the mask rebuilding ratio.
 The number of mask rebuilding pairs $N$ is calculated as:
@@
 The subscript $i$ denotes the number of values that exceeds zero, indicating that the growing weight is more important than the pruning weight.
+III-C Procedure
+Like most post-training pruning methods [15, 8], LLM-Barber is executed within a single global LLM forward pass, with local backward passes for gradient computation in each block.
+Figure 2 illustrates the LLM-Barber workflow, which consists of four stages:
+(1) Sparsity Mask Initialization.
+Initialize a preliminary sparsity mask from the dense network by a post-training pruning technique.
+(2) Block-aware Reconstruction Error Computation.
+We use a block-aware reconstruction error to evaluate the discrepancy between the dense and sparse model outputs.
+(3) Back-propagation for Gradients.
+Gradients are automatically derived via back-propagation, and the product of weights and gradients serves as the pruning metric.
+(4) Sparsity Mask Rebuilding.
+Masks are sorted by pruning metric, unpruned weights in ascending order while pruned weights in descending order.
+We rebuild the weight masks by growing newly significant weights and pruning those became non-salient with a specific mask rebuilding ratio.
+Ultimately, we recalculated the block-aware reconstruction error to assess the enhancement brought by our LLM-Barber.
+Experimental results reveal a substantial decrease in error, markedly decelerating the rate of error accumulation.
+Input: Calibration samples $\mathbf{X}$, a block’s weights $\{\mathbf{W}^{l}\}^{L}_{l=1}$ and initial masks $\{\mathbf{M_{i}}^{l}\}^{L}_{l=1}$, mask rebuilding ratio $\alpha$.
+Output: Rebuilt sparsity masks $\{\mathbf{M_{r}}^{l}\}^{L}_{l=1}$.
+[…]
+[…]
+3:for $l$ in $\{1,2,\dots,L\}$ do
+4:  $\mathbf{M_{r}}^{l}\leftarrow\mathbf{M_{i}}^{l}$
+5:  $\mathbf{S}^{l}\leftarrow|\mathbf{W}^{l}|\cdot|\mathbf{G}^{l}|$
+[…]
+7:  for $n$ in $\{1,2,\dots,N\}$ do
+8:   Obtain growing index $gi,gj$ via Eq.(10)
+9:   Obtain pruning index $pi,pj$ via Eq.(11)
+10:   $\mathbf{M_{r}}^{l}_{pi,pj}=1$triangleright Weight Growing.
+11:   $\mathbf{M_{r}}^{l}_{gi,gj}=0$triangleright Weight Pruning.
+12:  end for
+13:end for
+[…]
+15: Identify improvement by comparing $\mathbf{E_{i}}$, $\mathbf{E_{r}}$.
+16:return rebuilt sparsity masks $\{\mathbf{M_{r}}^{l}\}^{L}_{l=1}$.
+Algorithm 1 Pseudocode of LLM-Barber.
+III-D Structured N:M Sparsity
+While LLM-Barber primarily targets unstructured sparsity, it can be adapted for structured N:M sparsity.
+Groups of M weights are pruned to retain only N non-zero weights.
+During mask rebuilding, it divides each M-group into N pairs (one pruned and one non-pruned weight), then sorts them by output channel to identify mask rebuilding pairs.
+This method optimizes sparsity mask, leveraging N:M sparsity while maintaining model performance.
+TABLE I: WikiText-2 perplexity comparison at 50% sparsity rate.
+Bold results show improvements of integrating LLM-Barber.
+Underscored results indicate best performance in each LLM.WR represents weight reconstruction.
 LLaMA1 LLaMA2 LLaMA3 OPT
@@
 w/ LLM-Barber 7.118 6.091 6.868 5.918 9.451 11.95 11.71
-Table 1: WikiText-2 perplexity comparison for pruning LLM models at 50% sparsity rate.
-Bold results show improvements of integrating LLM-Barber.
-Underscored results indicate best performance in each LLM.
-WR represents weight reconstruction.
-Input: Calibration samples $\mathbf{X}$, a block’s weights $\{\mathbf{W}^{l}\}^{L}_{l=1}$ and initial masks $\{\mathbf{M_{i}}^{l}\}^{L}_{l=1}$, mask rebuilding ratio $\alpha$.
-Output: Rebuilt sparsity masks $\{\mathbf{M_{r}}^{l}\}^{L}_{l=1}$.
-1:  $\mathbf{E_{i}}\leftarrow\text{Block}\mathbf{(W,X)}-\text{Block}\mathbf{(W\odot M_{i},X)}$
-2:  $\{\mathbf{G}^{l}\}^{L}_{l=1}\leftarrow$ backpropagation for gradients via $\mathbf{E_{i}}$.
-3:  for $l$ in $\{1,2,\dots,L\}$ do
-4:     $\mathbf{M_{r}}^{l}\leftarrow\mathbf{M_{i}}^{l}$triangleright Initialize rebuilt sparsity mask.
-5:     $\mathbf{S}^{l}\leftarrow|\mathbf{W}^{l}|\cdot|\mathbf{G}^{l}|$triangleright Obtain importance score.
-[…]
-7:     for $n$ in $\{1,2,\dots,N\}$ do
-8:        Obtain growing index $gi,gj$ via Eq.(10)
-9:        Obtain pruning index $pi,pj$ via Eq.(11)
-10:        $\mathbf{M_{r}}^{l}_{pi,pj}=1$triangleright Weight Growing.
-11:        $\mathbf{M_{r}}^{l}_{gi,gj}=0$triangleright Weight Pruning.
-12:     end for
-13:  end for
-14:  $\mathbf{E_{r}}\leftarrow\text{Block}\mathbf{(W,X)}-\text{Block}\mathbf{(W\odot M_{r},X)}$
-15:  Identify improvement by comparing error $\mathbf{E_{i}}$ and $\mathbf{E_{r}}$.
-16:  return rebuilt sparsity masks $\{\mathbf{M_{r}}^{l}\}^{L}_{l=1}$.
-Algorithm 1 Pseudocode of LLM-Barber.
-Procedure
-Like most post-training pruning methods (Han et al.
-2015; Sun et al.
-2023), LLM-Barber is executed within a single global LLM forward pass, with local backward passes for gradient computation in each block.
-Figure 2 illustrates the LLM-Barber workflow, which consists of four stages:
-  • Sparsity Mask Initialization. We apply a post-training pruning technique to initialize a preliminary sparsity mask from the dense network.
-  • Block-aware Reconstruction Error Computation. We use a block-aware reconstruction error to evaluate the discrepancy between the dense and sparse model outputs.
-  • Back-propagation for Gradients. Gradients are automatically derived via back-propagation, and the product of weights and gradients serves as the pruning metric.
-  • Sparsity Mask Rebuilding.
-Masks are sorted based on the pruning metric, unpruned weights in ascending order while pruned weights in descending order.
-We rebuild the weight masks by growing newly significant weights and pruning those became non-salient with a certain percentage named mask rebuilding ratio.
-Ultimately, we recalculated the block-aware reconstruction error to assess the enhancement brought by our LLM-Barber.
-Experimental results reveal a substantial decrease in error, markedly decelerating the rate of error accumulation.
-Structured N:M Sparsity
-While LLM-Barber primarily targets unstructured sparsity, it can be adapted for structured N:M sparsity.
-Here, groups of M weights are pruned to retain only N non-zero weights.
-During mask rebuilding, LLM-Barber divides each M-group into N pairs (each pair contains one pruned and one non-pruned weight), then sorts these pairs by output channel to determine which are mask rebuilding pairs.
-This method creates an optimal sparse mask, leveraging N:M sparsity while maintaining model performance.
-Experiment
-Experiment Settings
+IV Experiment
+IV-A Experiment Settings
 Setup.
@@
 Models & Datasets.
-LLM-Barber is evaluated on the LLaMA family, including LLaMA-7B/13B (Touvron et al.
-2023a), LLaMA2-7B/13B (Touvron et al.
-2023b), and LLaMA3-8B (AI@Meta 2024), as well as the OPT model series: OPT-6.7B/13B (Zhang et al.
-2022).
+LLM-Barber is evaluated on the LLaMA family, including LLaMA-7B/13B [16], LLaMA2-7B/13B [17], and LLaMA3-8B [18], as well as the OPT model series: OPT-6.7B/13B [19].
 Notably, LLM-Barber is broadly applicable to any Transformer-based LLMs with Self-Attention and MLP blocks.
-Following previous works, we use 128 segments of 2048 tokens from the C4 dataset (Raffel et al.
-2020) for mask rebuilding.
+Following previous works, we use 128 segments of 2048 tokens from the C4 dataset [20] for mask rebuilding.
 Evaluation.
 To comprehensively assess LLM-Barber, we conduct rigorous evaluations on perplexity and zero-shot accuracy.
-Perplexity is measured on the validation sets of benchmarks such as WikiText-2 (Merity et al.
-2017), PTB (Marcus et al.
-1994), and C4 (Raffel et al.
-2020).
-Zero-shot accuracy is assessed using the EleutherAI LM Harness (Gao et al.
-2023) across six benchmarks: BoolQ (Clark et al.
-2019), RTE (Wang et al.
-2018), HellaSwag (Zellers et al.
-2019), ARC Easy and Challenge (Clark et al.
-2018), and OpenbookQA (Mihaylov et al.
-2018).
-Baselines. The results of LLM-Barber are compared with the following established post-training pruning methods:
-  • Magnitude Pruning (Han et al. 2015) eliminates weights based only on their magnitudes;
-  • SparseGPT (Frantar and Alistarh 2023) identifies weights importance by using second-order information;
-  • Wanda (Sun et al. 2023) determines weights to be pruned by the weight magnitude multiplied by input activation.
-Main Results
+Perplexity is measured on the validation sets of benchmarks such as WikiText-2 [21], PTB [22], and C4 [20].
+Zero-shot accuracy is assessed using the EleutherAI LM Harness [23] across six benchmarks: BoolQ [24], RTE [25], HellaSwag [26], ARC Easy and Challenge [27], and OpenbookQA [28].
+Baselines.
+The results of LLM-Barber are compared with the following established post-training pruning methods: (1) Magnitude Pruning [15] eliminates weights based only on their magnitudes; (2) SparseGPT [4] identifies weights importance by using second-order information; (3) Wanda [8] determines weights to be pruned by the weight magnitude multiplied by input activation.
+IV-B Main Results
 Unstructured Sparsity.
-LLM-Barber effectively prunes the LLaMA and OPT models, achieving 50% unstructured sparsity without requiring supplementary weight reconstruction, as detailed in Table 1.
+LLM-Barber effectively prunes the LLaMA and OPT models, achieving 50% unstructured sparsity without requiring supplementary weight reconstruction, as detailed in Table I.
 LLM-Barber demonstrates the capability to rebuild the sparsity masks initialized by other pruning methods in a single forward pass, significantly outperforming conventional pruning baselines.
@@
 Notably, LLM-Barber achieves robust improvements even with poorly performing initial sparsity masks, such as magnitude pruning, where it reduces perplexity from 205.5 to 10.98, which is an impressive and substantial enhancement.
+Varying Sparsity Levels.
+We conduct experiments on varying sparsity levels for unstructured pruning in LLaMA3-8B as shown in Table II.
+LLM-Barber consistently increases perplexity across all initialization methods, with magnitude pruning showing the most significant improvement.
+TABLE II: WikiText-2 perplexity performance for pruning LLaMA3-8B at varying sparsity rate.
 Sparsity 60% 70% 80% 90%
 Magnitude 3.39e4 1.62e6 8.55e7 2.26e7
-w/ LLM-Barber 28.14 2.08e2 1.09e3 7.55e4
+w/ Barber 28.14 2.08e2 1.09e3 7.55e4
 Wanda 23.57 1.28e2 8.54e2 1.28e4
-w/ LLM-Barber 22.04 1.07e2 5.70e2 7.07e3
-Table 2: WikiText-2 perplexity performance for pruning LLaMA3-8B at varying sparsity rate.
-Varying Sparsity Levels.
-We conduct experiments on varying sparsity levels for unstructured pruning in LLaMA3-8B as shown in Table 2.
-LLM-Barber consistently increases perplexity across all initialization methods, with magnitude pruning showing the most significant improvement.
+w/ Barber 22.04 1.07e2 5.70e2 7.07e3
+Structured N:M Sparsity.
+In contrast to unstructured sparsity, employing N:M fine-grained sparsity can provide more tangible acceleration benefits when leveraging NVIDIA Ampere’s sparse tensor cores [29].
+Therefore, we also evaluate the effectiveness of our LLM-Barber in partial LLaMA models on the N:M fine-grained sparsity pattern as shown in Table III.
+TABLE III: WikiText-2 perplexity comparison for pruning LLaMA family with structured N:M pattern.
 Method Sparsity V1-7B V1-13B V2-7B
 Magnitude 4:8 16.83 13.72 15.91
-w/ LLM-Barber 4:8 8.852 7.137 9.345
+w/ Barber 4:8 8.852 7.137 9.345
 SparseGPT 4:8 8.608 7.437 8.495
-w/ LLM-Barber 4:8 8.191 7.085 8.003
+w/ Barber 4:8 8.191 7.085 8.003
 Magnitude 2:4 42.56 18.32 37.76
-w/ LLM-Barber 2:4 11.04 9.006 13.47
+w/ Barber 2:4 11.04 9.006 13.47
 SparseGPT 2:4 11.55 9.116 10.94
-w/ LLM-Barber 2:4 10.14 8.517 9.806
-Table 3: WikiText-2 perplexity comparison for pruning LLaMA family with structured N:M pattern.
-Structured N:M Sparsity.
-In contrast to unstructured sparsity, employing N:M fine-grained sparsity can provide more tangible acceleration benefits when leveraging NVIDIA Ampere’s sparse tensor cores (Choquette et al.
-2021).
+w/ Barber 2:4 10.14 8.517 9.806
+TABLE IV: Zero-shot performance comparison of LLaMA series on six tasks at 50% sparsity rate.
+Bold results show improvements of integrating LLM-Barber.
+Underscored results indicate the best performance for each tasks.
 Model Method BoolQ RTE HellaSwag ARC-e ARC-c OBQA Mean
@@
 w/LLM-Barber 78.59 61.37 51.01 71.46 39.85 29.00 55.21
-Table 4: Zero-shot performance comparison for pruning LLaMA series on six evaluated tasks at 50% sparsity rate.
-Bold results show improvements of integrating LLM-Barber.
-Underscored results indicate the best performance in each zero-shot tasks.
-Therefore, we also evaluate the effectiveness of our LLM-Barber in partial LLaMA models on the N:M fine-grained sparsity pattern as shown in Table 3.
 Zero-shot Performance.
-Following previous works (Frantar and Alistarh 2023; Sun et al.
-2023), we evaluated the LLaMA models on six diverse zero-shot tasks.
-The results are summarized in Table 4, where models are pruned to unstructured 50% sparsity.
+Following previous works [4, 8], we evaluated the LLaMA models on six diverse zero-shot tasks.
+The results are summarized in Table IV, where models are pruned to unstructured 50% sparsity.
 Averaging the accuracy across the six evaluated tasks, it becomes apparent that LLM-Barber possesses the capability to identify a more effective network than those obtained via the initialization methods.
@@
 However, it is worth noting that there is no single universally superior performer for the remaining tasks in the evaluation set, with the initial pruning methods sometimes matching or even marginally exceeding the results obtained through LLM-Barber.
-Mask Rebuilding Ratio Selection
-Figure 3: Relationship between importance score distribution of mask rebuilding pair and WikiText-2 perplexity results at varying pruning granularities in LLaMA-7B.
-The green dashed line indicates the best choice of mask rebuilding ratio.
+TABLE V: Comparison of DSnoT vs.
+LLM-Barber on LLaMA3-8B at 50% unstructured sparsity.
+We report WikiText-2 perplexity (lower is better) and zero-shot accuracy (%) on five downstream tasks.
+Bold indicates best result for each setting.
+Method WikiText-2 ARC-E ARC-C BoolQ OBQA RTE
+Dense (no pruning) 6.136 75.29 41.89 75.08 34.40 66.79
+Magnitude (50%) 205.5 46.59 25.09 42.87 22.00 53.07
+    w/ DSnoT 47.52 60.52 32.16 49.66 25.20 54.87
+    w/ LLM-Barber 10.98 72.05 37.49 72.72 27.40 54.87
+Wanda (50%) 9.821 71.38 40.10 78.41 29.40 60.29
+    w/ DSnoT 9.562 71.25 40.01 75.99 29.00 61.81
+    w/ LLM-Barber 9.451 71.46 39.85 78.59 29.00 61.37
+Comparison with DSnoT.
+We further compare LLM-Barber with the recent training-free pruning method DSnoT [7] to highlight their differences in approach and performance.
+DSnoT also aims to improve sparse LLMs without weight retraining, but it does so by iteratively pruning and regrowing weights in multiple cycles to minimize reconstruction In contrast, LLM-Barber performs all mask rebuilding in a single global forward pass (with localized backward gradient computations per block), making it a true one-shot method.
+This fundamental distinction gives LLM-Barber an efficiency edge: it avoids the repeated passes of DSnoT, enabling faster one-shot deployment.
+Moreover, DSnoT imposes a fixed number of iterations (a maximum cycle count) for mask update in each layer, which can lead to sub-optimal results when the initial mask is poor.
+LLM-Barber is not bound by a preset iteration limit; instead, it dynamically rebuilds the mask based on our weight-importance metric until all significant weights are retained.
+This adaptability allows LLM-Barber to recover much more performance, especially starting from an inferior initial mask.
+Importantly, LLM-Barber’s block-aware strategy provides greater flexibility via global coordination across Transformer blocks, whereas DSnoT treats layers more independently (resembling layerwise fine-tuning).
+By coordinating mask updates at the block level, LLM-Barber better preserves end-to-end model quality.
+In summary, qualitatively LLM-Barber offers higher efficiency, flexibility, and overall performance than DSnoT.
+Its one-shot block-aware mask rebuilding avoids DSnoT’s iterative overhead and limitation, yielding a more globally coordinated pruning outcome.
+These advantages make LLM-Barber well-suited for scenarios that require rapid one-shot compression for LLM deployment without sacrificing accuracy.
+[…]
+IV-C Mask Rebuilding Ratio Selection
+Figure 3: The importance score distribution of mask rebuilding pairs and WikiText-2 perplexity results at varying pruning granularities in LLaMA-7B, with the green dashed line marking the optimal mask rebuilding ratio.
 A critical aspect of the LLM-Barber method lies in determining the optimal mask rebuilding ratio to achieve peak accuracy.
@@
 Fortunately, this distribution often exhibits a distinct pattern of outliers, facilitating rapid identification of an appropriate mask rebuilding ratio.
-We have plotted the score distribution of mask rebuilding pairs for various pruning granularity and initialization methods on LLaMA-7B, as well as perplexity performance corresponding to mask rebuilding ratios in Figure 3.
+We plotted the score distribution of mask rebuilding pairs across various pruning granularity and initialization methods on LLaMA-7B, along with perplexity performance corresponding to mask rebuilding ratios (Figure 3).
 The results reveal a strong correlation between the distribution of outliers in mask rebuilding pairs and the optimal mask rebuilding ratio.
 For instance, outliers are significantly distributed within the top 10% with Magnitude pruning, thus selecting a 10% mask rebuilding ratio yields the overall optimal solution.
-Similarly, with the initialization method Wanda, a significant outlier distribution around the 1% mark corresponds to optimal results near a 1% mask rebuilding ratio.
-Thus, LLM-Barber can preemptively narrow the search range for the optimal mask rebuilding ratio by analyzing outlier distributions.
-This approach allows LLM-Barber to adapt flexibly to various reconstruction masks without extensive searching.
+Similarly, the Wanda initialization method shows a notable outlier distribution around the 1% mark corresponds to optimal results near a 1% mask rebuilding ratio.
+Thus, LLM-Barber can preemptively narrow the search range for the optimal mask rebuilding ratio by analyzing outlier distributions, allowing for flexible adaptation to various reconstruction masks without extensive searching.
 It is worth noting that LLM-Barber identifies a larger proportion of outliers in less effective initial masks, which corresponds to a more substantial rebuilding of the mask.
 This is why LLM-Barber provides a more significant improvement for masks with poorer initial sparsity mask with a more aggressive mask rebuilding strategy.
-Ablation Study
-Given the outstanding potential of LLM-Barber, we have comprehensively analyzed the impact of three critical factors to assess the robustness and effectiveness of our LLM-Barber method: pruning granularity, pruning metric, and calibration data size.
-These factors were chosen to gain a deeper understanding of how different configurations influence the performance of pruned models and to demonstrate LLM-Barber’s versatility across various settings.
+TABLE VI: Pruning granularity ablation in LLaMA3-8B. Bold results show best granularity of each row.
+Method Pruning Granularity
+w/ Barber Block Layer Input Output
+Magnitude 10.98 11.06 72.48 11.81
+SparseGPT 9.380 9.348 9.418 9.567
+Wanda 9.626 9.633 9.849 9.451
+IV-D Ablation Study
+Given the significant potential of LLM-Barber, we analyzed three critical factors to assess the robustness and effectiveness of our method: pruning granularity, pruning metric, and calibration data size.
+These factors were chosen to gain a deeper understanding of how different configurations affect the performance of pruned models and to demonstrate LLM-Barber’s versatility across various settings.
+TABLE VII: Ablation of pruning metric in LLaMA3-8B. Bold results support the pruning metric $|\mathbf{W}||\mathbf{\partial E/\partial W}|$.
+Method Pruning Metric
+w/ Barber $||\mathbf{W}|$$||\mathbf{\partial E/\partial W}|$$|\mathbf{W}||\mathbf{\partial E/\partial W}|$
+Magnitude 186.5 14.43 10.98
+SparseGPT 9.544 9.457 9.348
+Wanda 9.880 9.699 9.451
 Pruning Granularity.
 LLM-Barber dynamically selects different pruning granularities to adapt to varying initialization methods.
-In this paper, we evaluated the impact of four levels of granularities: block-wise, layer-wise, input-wise, and output-wise pruning, as shown in Table 5.
+In this paper, we evaluated the impact of four levels of granularities: block-wise, layer-wise, input-wise, and output-wise pruning, as shown in Table VI.
 For magnitude pruning, the block-wise granularity yields the best performance, while the output-wise granularity delivers the lowest perplexity for Wanda pruning.
 By tailoring the pruning granularity to the particular pruning approach, LLM-Barber can consistently achieve optimal model performance.
-Method Pruning Granularity
-w/ LLM-Barber Block Layer Input Output
-Magnitude (205.5) 10.98 11.06 72.48 11.81
-SparseGPT (9.994) 9.380 9.348 9.418 9.567
-Wanda (9.821) 9.626 9.633 9.849 9.451
-Table 5: Ablation of pruning granularity in LLaMA3-8B. Perplexity in brackets show the WikiText-2 score without LLM-Barber. Bold results show best granularity of this row.
+Figure 4: Ablation of calibration size in LLaMA3-8B. LLM-Barber is robust across varying calibration size.
+TABLE VIII: Inference performance of different quantization methods on FPGA platform.
+Results are reported for Qwen2-7B on EdgeLLM with the ARC-Challenge, ARC-Easy, and BoolQ tasks in zero-shot setting.
+Method ARC-Challenge ARC-Easy BoolQ
+TPS $\uparrow$ Latency(ms) $\downarrow$ Acc(%) $\uparrow$ TPS $\uparrow$ Latency(ms) $\downarrow$ Acc(%) $\uparrow$ TPS $\uparrow$ Latency(ms) $\downarrow$ Acc(%) $\uparrow$
+GPTQ 30.01 1.98 46.28 29.88 1.75 64.71 30.31 0.65 56.88
+AWQ 28.13 2.23 35.38 28.85 2.02 57.62 28.47 1.98 51.84
+APTQ+Ours 30.37 1.92 56.28 30.14 1.58 67.16 31.48 0.48 60.94
 Pruning Metric.
-We analyze the effect of different pruning metrics, focusing on weight magnitude, gradient magnitude, and the product of weight and gradient.
-As seen in Table 6, the product of weight and gradient consistently outperforms other metrics, achieving the lowest perplexity of 10.98 for Magnitude, 9.348 for SparseGPT, and 9.451 for Wanda.
+We analyze the effect of different pruning metrics, weight magnitude, gradient magnitude, and product of weight and gradient.
+As shown in Table VII, the product of weight and gradient consistently outperforms the others, achieving the lowest perplexity of 10.98 for Magnitude, 9.348 for SparseGPT, and 9.451 for Wanda.
 This confirms the effectiveness of our product-based pruning metric across various methods.
-Method Pruning Metric
-w/ LLM-Barber $|\mathbf{W}|$ $|\mathbf{\partial E/\partial W}|$ $|\mathbf{W}|\cdot|\mathbf{\partial E/\partial W}|$
-Magnitude (205.5) 186.5 14.43 10.98
-SparseGPT (9.994) 9.544 9.457 9.348
-Wanda (9.821) 9.880 9.699 9.451
-Table 6: Ablation of pruning metric in LLaMA3-8B. Bold results support the pruning metric $|\mathbf{W}|\cdot|\mathbf{\partial E/\partial W}|$.
 Calibration Data Size.
@@
 LLM-Barber outperforms SparseGPT even with just a single sample, underscoring its robustness across different sample sizes.
-Figure 4: Ablation of calibration data size in LLaMA3-8B. LLM-Barber is robust across varying calibration data size.
-Conclusion
-We propose LLM-Barber (Block-Aware Rebuilder for Sparsity Mask in One-Shot), a novel framework that rebuilds sparsity mask to optimize LLM post-training pruning.
-By integrating a block-aware approach across Self-Attention and MLP blocks, LLM-Barber effectively reallocates sparsity to improve accuracy without the need for extensive fine-tuning.
-Specifically, LLM-Barber identifies novel importance score after mask initialization and rebuilds the sparsity mask with mask rebuilding pairs, simultaneously applying new sparsity masks to weights that have become less critical, thereby optimizing overall model performance.
-By utilizing the novel pruning metric as the product of weights and gradients, our approach enables precise and efficient reallocation of sparsity mask.
-Extensive experiments on pruning LLaMA series models demonstrate that LLM-Barber achieves state-of-the-art results in both perplexity and zero-shot performance within the domain of post-training pruning.
+IV-E Deployment-Aware Quantization Compatibility
+TABLE IX: Comparative Throughput Analysis of Quantization Methods Across GPU Frameworks with LLM-Barber Optimization
+Method Throughput(Tokens/sec) Baseline
+CUDA Triton Marlin ExLlama BF16
+GPTQ 37.38 15.90 38.03 50.59 44.30
+AWQ 38.10 16.21 43.46 51.63 44.30
+APTQ+Ours 38.29 16.42 47.13 52.29 44.30
+LLM-Barber’s block-aware sparsity method provides an efficient compression solution when combined with post-training quantization like APTQ [12].
+This integration enhances both model size reduction and inference speed.
+In this section, we present experimental results validating the compatibility and acceleration benefits of combining LLM-Barber with APTQ.
+We evaluate the model’s performance on both GPU and FPGA platforms, comparing tokens per second (TPS), latency, and accuracy across different quantization methods: GPTQ [11], AWQ [30], and APTQ.
+As shown in Table IX, LLM-Barber combined with APTQ achieves higher TPS and lower latency on both platforms, while maintaining competitive accuracy.
+Additionally, we extend this analysis to FPGA hardware using the EdgeLLM [31] platform, where we compare the performance of the quantization methods on the ARC-Challenge, ARC-Easy, and BoolQ datasets.
+As detailed in Table VIII, APTQ, when integrated with LLM-Barber’s block-sparsity, delivers the highest TPS, lowest latency, and best accuracy across all tasks.
+This validates the effectiveness of LLM-Barber’s block-aware sparsity and APTQ’s quantization on FPGA, demonstrating that our method significantly accelerates deployment while preserving model performance.
+In summary, the combination of LLM-Barber’s block-sparsity approach with APTQ optimizes both inference speed and deployment flexibility.
+Our results confirm that this joint optimization provides higher TPS and lower latency on GPU while achieving substantial acceleration on FPGA platforms.
+These findings highlight the potential of integrating sparse pruning with quantization for end-to-end deployment of LLMs, enabling scalable, efficient, and hardware-agnostic solutions across different platforms, such as GPU and FPGA.
+Future work will explore further quantization schemes and optimizations for edge deployment.
+We envision expanding this approach to larger model architectures and more diverse hardware environments, making large-scale LLM deployment more efficient and accessible.
+V Conclusion
+We introduced LLM-Barber, a one-shot pruning framework that achieves aggressive compression of large language models without retraining.
+By leveraging a block-aware mask rebuilding strategy and a weight-gradient importance metric, LLM-Barber removes 50–60% of weights while preserving state-of-the-art accuracy.
+Our approach prunes multi-billion-parameter models in minutes and incorporates hardware-aware optimizations, such as structured block sparsity and low-bit quantization, to ensure effective deployment on modern accelerators.
+Experimental results demonstrate that LLM-Barber outperforms unstructured pruning baselines like SparseGPT and Wanda in both perplexity and downstream task accuracy.
+Additionally, the FPGA deployment of a pruned+quantized model achieves significant speedups and energy savings, proving LLM-Barber’s hardware efficiency.
+In conclusion, LLM-Barber bridges the gap between theoretical model compression and practical deployment, enabling efficient LLM inference in resource-constrained environments.
+Future work will scale this approach to larger models and integrate it with other compression techniques.
+Acknowledgement
+This work was supported by the Shenzhen Science and Technology Program under Grant No.
+KQTD20200820113051096, Shenzhen Science and Technology Program under Grant No.
+JCYJ20220818100217038, the National Natural Science Foundation of China under Key Program Grant No.
+62034007, the Science and Technology Innovation Key R&D Program of Chongqing under Grant No.
+CSTB2023TIAD-STX0001, and the Hong Kong Theme-based Research Scheme (TRS) project T45-701/22-R of the Research Grants Council (RGC).